Solve the equation :


(7y-3x+3) dy +(3y-7x+7) dx=0

1
Expert's answer
2022-01-31T18:01:47-0500

Solution:

"(7y-3x+3)dy + (3y-7x+7)dx =0

\\\Rightarrow (3x-7y-3)dy=(3y-7x+7)dx"

"\Rightarrow\dfrac{dy}{dx}=\dfrac{3y-7x+7}{3x-7y-3}\\
now,let ~x=x'+h~and~y=y'+k\\
\Rightarrow \dfrac{dy'}{dx'}=\dfrac{3y'-7x'+3k-7h+7}{3x'-7y'+3h-7k-3}\\
putting~h=1~and~k=0,we~get\\
\Rightarrow\dfrac{dy'}{dx'}=\dfrac{3y'-7x'}{3x'-7y'}\\
now,let~ y'=vx'~so~that\\
\Rightarrow\dfrac{dy'}{dx'}=v+x'\dfrac{dv}{dx'}\\
we~have\\
\Rightarrow v+x'\dfrac{dv}{dx'}=\dfrac{3vx'-7x'}{3x'-7vx'}\\
\Rightarrow x'\dfrac{dv}{dx'}=\dfrac{3v-7}{3-7v}-v\\
\Rightarrow x'\dfrac{dv}{dx'}=\dfrac{3v-7-3v+7v^2}{3-7v}\\
\Rightarrow x'\dfrac{dv}{dx'}=\dfrac{7v^2-7}{3-7v}\\
\Rightarrow\dfrac{dx'}{x'}=\dfrac{3-7v}{7v^2-7}dv\\
integrating~both~sides\\
\Rightarrow\int \dfrac{dx'}{x'}=\int \dfrac{3}{7v^2-7}dv-\int \dfrac{7v}{7v^2-7}dv\\
\Rightarrow ln|x'|+c= \dfrac{3}{14}ln \dfrac{|v-1|}{|v+1|}-\dfrac{1}{2}ln|{v^2-1}|\\
\Rightarrow ln|x'|+c= \dfrac{3}{14}ln \dfrac{|y'-x'|}{|y'+x'|}-\dfrac{1}{2} ln|\dfrac{(y')^2-(x')^2}{(x')^2}|\\
Recall~v=\dfrac{y'}{x'}\\
\Rightarrow ln|x'|+c= \dfrac{3}{14}ln {|y'-x'|}- \dfrac{3}{14}{|y'+x'|}-\dfrac{1}{2} ln|{y'-x'}|+\dfrac{1}{2} ln|{y'+x'}+\dfrac{1}{2} ln|({x'})^2|\\
\Rightarrow ln|x'|+c= -\dfrac{2}{7}ln {|y'-x'|}-\dfrac{5}{7}ln {|y'+x'|}+\dfrac{1}{2} ln|({x'})^2|\\
\Rightarrow 2ln|y'-x'|+5ln|y'+x'|=C\\
\Rightarrow2ln|y-x+1|+5ln|y+x-1|=C\\ 

(Since~y'=y~and~x'=(x-1))"